# Text-to-Image Rectified Flow as Plug-and-Play Priors

### v2

@@ A.2 Additional Related Works
Diffusion as plug-and-play priors. Our work is greatly motivated by the development of diffusion based priors. The earliest work of diffusion prior [13] requires to backpropagate through the diffusion U-Net. Dreamfusion [14] recognizes that such a backpropagate process greatly hurt the performance of diffusion priors. By ignoring the U-Net Jacobian, the SDS loss proposed in Dreamfusion can be used for 3D generation and image editing. However, the initial version of the SDS loss suffers greatly from the lack of details and diversity. DDS loss [17] proposes an improved version of the SDS to improve image editing by taking the difference between the current SDS and the source image SDS. VSD loss improves the SDS loss for 3D generation problems. Specifically, it first trains a LoRA of the current 3D model and then takes the difference between the diffusion SDS and LoRA SDS. Although lots of methods have been proposed for diffusion models, this is the first work that studies how to effectively use rectified flow as priors.
A.3 Algorithm for RFDS and iRFDS
1 Initialize the learnable parameter $\theta$
2
3while Not Converge do
@@ 5       Sample random noise $\epsilon$
7      Optimize $\theta$ with $\epsilon$ based on RFDS (Eq. 6)
8
9
RETURN $\theta$
@@
1 Initialize the learnable parameter $\epsilon$
2 Get initial image $x$
3
4while Not Converge do
@@ 5       Sample random timestep $t$
7      Optimize $\epsilon$ using fixed $x$ based on iRFDS (Eq. 8)
8
9
RETURN $\epsilon$
@@
A.4 Additional Ablation Experiments
Ablation Results on CFG. The scale of classifier-free guidance (CFG) [25] plays a crucial role in diffusion-based methods, such as SDS and VSD. We observe a very similar phenomenon with the rectified flow priors. As demonstrated in Fig. 9, both of the proposed methods require a CFG greater than 10 to learn reasonable shapes. However, when the CFG becomes excessively large, the 3D objects generated by RFDS exhibit over-saturated colors. In contrast, RFDS-Rev remains robust to large CFG values, even when the CFG exceeds 2000.
Figure 9: Ablation experiments of classifer free guidance scale on text-to-3D generation. Prompt: A DSLR image of a hamburger. Top: RFDS with different CFG. Bottom: RFDS-Rev.
RFDS-Rev vs. RFDS-VSD. As mentioned in the main text, some of the existing methods aimed at improving the diffusion models can be used directly on rectified flow based methods. We explore to combine VSD with the baseline RFDS, denoted as RFDS-VSD. Specifically, we train a rectified flow LoRA model based on the current rendered images and then calculate the gradients by taking the difference between RFDS and RFDS-LoRA following the VSD setting. Results are shown in Fig. 10. We observe that the RFDS-Rev generates much better results compared with RFDS-VSD, even though the VSD version requires more computations.
@@ A.4 Additional Ablation Experiments
A.5 Comparison of Convergence Speed with Diffusion-based Methods.
@@
A.6 Comparison of the Computational Cost.
We evaluate the computational costs by examining the number of forward and backward passes of the diffusion or rectified flow network required in 1 optimization iteration. Results are list in Table. 3. The RFDS baseline has the same computational demands as the SDS loss. Due to the calculation of CFG [25], they both require two forward passes. RFDS-Rev requires only one additional forward pass, whereas VSD needs two additional forward passes and one additional costly backward pass.
@@
Method SDS [14] VSD [15] DDS [17] RFDS iRFDS RFDS-Rev
@@
A.7 More results of image inversion and editing using iRFDS.
@@
A.8 More results of text-to-3D generation
We show more qualitative results of text-to-3D generation in Fig. 13 and Fig. 14.
Figure 13: More results on text-to-3D generation
Figure 14: More results on text-to-3D generation
A.9 iRFDS user study details
@@ A.8 More results of text-to-3D generation
Figure 15: User study screenshot

### v3

@@ C Additional Related Works
Diffusion as plug-and-play priors. Our work is greatly motivated by the development of diffusion-based priors. The earliest work of diffusion prior (Graikos et al., 2022) requires to backpropagate through the diffusion U-Net. Dreamfusion (Poole et al., 2022) recognizes that such a backpropagate process greatly hurts the performance of diffusion priors. By ignoring the U-Net Jacobian, the SDS loss proposed in Dreamfusion can be used for 3D generation and image editing. However, the initial version of the SDS loss suffers greatly from the lack of details and diversity. DDS loss (Hertz et al., 2023) proposes an improved version of the SDS to improve image editing by taking the difference between the current SDS and the source image SDS. VSD loss (Wang et al., 2023) improves the SDS loss for 3D generation problems. Specifically, it first trains a LoRA of the current 3D model and then takes the difference between the diffusion SDS and LoRA SDS. Although lots of methods have been proposed for diffusion models, this is the first work that studies how to effectively use rectified flow as priors.
D Algorithm for RFDS and iRFDS
We list the detailed algorithm for RFDS and iRFDS in Algorithm 2 and Algorithm 3.
1 Initialize the learnable parameter $\bm{\theta}$
2 while Not Converge do
@@ 4       Sample random noise $\epsilon$
5       Optimize $\bm{\theta}$ with $\bm{\epsilon}$ based on RFDS (Eq. 8)
@@
RETURN $\bm{\theta}$
@@
1 Initialize the learnable parameter $\bm{\epsilon}$
2 Get initial image $\bm{x}$
3 while Not Converge do
@@ 4       Sample random timestep $t$
5       Optimize $\bm{\epsilon}$ using fixed $\bm{x}$ based on iRFDS (Eq. 10)
@@
RETURN $\bm{\epsilon}$
@@
E Additional Ablation Experiments
Ablation Results on CFG. The scale of classifier-free guidance (CFG) (Ho & Salimans, 2022) plays a crucial role in diffusion-based methods, such as SDS and VSD. We observe a very similar phenomenon with the rectified flow priors. As demonstrated in Fig. 8, both of the proposed methods require a CFG greater than 10 to learn reasonable shapes. However, when the CFG becomes excessively large, the 3D objects generated by RFDS exhibit over-saturated colors. In contrast, RFDS-Rev remains robust to large CFG values, even when the CFG exceeds 2000.
Figure 8: Ablation experiments of classifer free guidance scale on text-to-3D generation. Prompt: A DSLR image of a hamburger.
RFDS-Rev vs. RFDS-VSD. As mentioned in the main text, some of the existing methods aimed at improving the diffusion models can be used directly on rectified flow based methods. We explore to combine VSD with the baseline RFDS, denoted as RFDS-VSD. Specifically, we train a rectified flow LoRA model based on the current rendered images and then calculate the gradients by taking the difference between RFDS and RFDS-LoRA following the VSD setting. Results are shown in Fig. 9. Our experiments, conducted using the InstaFlow backbone, demonstrate that RFDS-Rev produces significantly better results compared to RFDS-VSD, despite RFDS-VSD requiring more computational resources. Notably, implementing VSD on the SD3 model presents significant challenges for most currently available commercial GPUs due to its requirement for fine-tuning the base model, which demands excessive GPU memory.
@@ E Additional Ablation Experiments
F Comparison of Convergence Speed.
@@
G Comparison of the Computational Cost.
We evaluate the computational costs by examining the number of forward and backward passes of the diffusion or rectified flow network required in 1 optimization iteration. Results are list in Table. 3. The RFDS baseline has the same computational demands as the SDS loss. Due to the calculation of CFG (Ho & Salimans, 2022), they both require two forward passes. RFDS-Rev requires only one additional forward pass, whereas VSD needs two additional forward passes and one additional costly backward pass.
@@
Method SDS Poole et al. (2022) VSD Wang et al. (2023) DDS Hertz et al. (2023) RFDS iRFDS RFDS-Rev
@@
H More results of image inversion and editing using iRFDS.
@@
I More results of text-to-3D generation
We show more qualitative results of text-to-3D generation in Fig. 12,Fig. 13, Fig. 14 and Fig. 15.
Figure 12: More results on text-to-3D generation. Model:SD3
Figure 13: More results on text-to-3D generation. Model:SD3
Figure 14: More results on text-to-3D generation. Model:InstaFlow
Figure 15: More results on text-to-3D generation. Model:InstaFlow
J iRFDS user study details
@@ I More results of text-to-3D generation
Figure 16: User study page